OpenStudy (anonymous):

If 8:x = 3:y then 8+x:x = ?:y would it b 3+y or just 3

OpenStudy (rulnick):

3+y

hero (hero):

Test it out using real numbers

OpenStudy (rulnick):

Assuming you mean (8+x):x = ?:y, then ? is (3+y).

OpenStudy (rulnick):

In fraction form, (8+x)/x = 1+(8/x) = 1 + 3/y = (3+y)/y.